OpenStudy (anonymous):

Factor the following trinomial. 64x^2+48xy+9y^2

OpenStudy (anonymous):

\[8^{2} \times x^{2} + 8 \times 6 \times xy + 3^{2}y^{2}\] This gives us : \[(8x + 3y)²\]